# Federated Learning for Smart Grid: A Survey on Applications and Potential Vulnerabilities

### v2

@@ 5.2.1. Detection-Filtering Methods
Detection-filtering methods aim to maximize robustness against attacks by identifying and removing malicious local model updates from the aggregation process. These methods typically rely on one or more statistical metrics (e.g., norms, cosine similarity, Euclidean distance) to assess local model updates. By comparing the extracted metrics with each update, they filter out updates that exhibit abnormal metric values, preventing them from influencing the global model. For example, Krum (krum) selects the most reliable local model that has the smallest sum of squared Euclidean distances to all other models as the output. Multi-Krum (krum) extends this by selecting multiple reliable models for averaging instead of just one. In (xu2022byzantine), the authors applied the KL-divergence to measure the distance between the empirical sign distributions of the local models in two consecutive rounds and used it to employ a weighted average of models for each client. The Divide-and-Conquer (DnC) method proposed in (dnc) calculates the product of a model with its principle component and removes the models with large projection values. FLTrust (fltrust) leverages a root dataset and trains a root global model based on it. After collecting local updates, FLTrust aggregates those updates with high Cosine similarity to root model updates. SignGuard (signguard) combines direction-based clustering and magnitude-based filtering to identify malicious model updates, where the direction and magnitude are obtained by the sign statistics and $L_{2}$-norm of the model update, respectively. Recently, a defense method called LASA that achieves SOTA results has been proposed in (LASA). It employs a layer-wise adaptive filter to adaptively select benign sparsified layers using both magnitude and direction metrics across all clients for aggregation.
5.2.2. Impact-Reduction Methods
@@
Impact-reduction methods try to integrate all model updates but employ strategies to reduce the impact of malicious updates. For instance, Trimmed mean (Trmean) proposed in (trmean) discards a certain percentage of the highest and lowest values among the received models from clients for each dimension. After this trimming, the mean of the remaining values is computed by the server, which mitigates the impact of extreme values on the aggregated model. Mhamdi et al. (bulyan) proposed Bulyan that combines Trmean and Multi-Krum to extract benign models. In (geomed), the authors proposed to use the geometric median (GeoMed) of all local models as the aggregated model, which is effective in filtering out the outlier values. RLR (rlr) uses the sign information of model updates to adjust the global learning rate for each model coordinate without distinguishing between benign and malicious updates, resulting in poor accuracy. Foolsgold (Foolsgold) assumes that the malicious updates are consistent with each other. It assigns aggregation weights to model updates based on the maximum Cosine similarity between the last layers of pairwise model updates. A higher Cosine similarity value indicates a higher probability that the updates are malicious, leading to smaller aggregation weights being assigned. SparseFed proposed in (sparsefed) sparsifies the aggregated model update at the server side, integrating with model clipping and error feedback, to mitigate the impact of malicious local model updates. In addition, FedVRDP proposed in (zhang2023byzantine) applies sparsification on local model updates using a mask obtained from the global model in the previous round and combines it with random Gaussian noise perturbation and variance reduction techniques, to achieve robustness and differential privacy protection simultaneously. A recent approach called Lockdown (lockdown) assumes that those malicious parameters served to recognize backdoor triggers will be deemed unimportant for benign clients and applies sparsification to prune unimportant parameters on the local model.
5.2.3. DP-based Methods
In DP for FL system, the goal is to ensure the privacy of individual client data during the model training process. In a system with $N$ clients, each client $n$ trains a local model on its own dataset $\mathcal{D}_{n}^{\text{train}}$ by minimizing a local loss function $f_{n}(\cdot)$. After training, the model parameters $\theta_{n}^{t}$ are updated as follows:
(9) $\theta_{n}^{t+1}=\theta_{n}^{t}-\eta\frac{1}{|\mathcal{D}_{n}^{\text{train}}|}\sum_{i=1}^{|\mathcal{D}_{n}^{\text{train}}|}\nabla_{\theta}f_{n}(x_{n,i},y_{n,i}).$
To preserve privacy, each client adds Gaussian noise $\mathcal{N}(0,(\sigma_{n}^{t})^{2}I)$ to the trained local model parameters:
(10) $\tilde{\theta}_{n}^{t+1}=\theta_{n}^{t+1}+\mathcal{N}(0,(\sigma_{n}^{t})^{2}I),$
where $I$ is the identity matrix, and $\sigma_{n}^{t}$ is the noise level at iteration $t$. After adding noise, clients upload their noised model parameters $\tilde{\theta}_{n}^{t+1}$ to the server, which aggregates them using federated averaging:
(11) $\theta_{G}^{t+1}=\frac{1}{N}\sum_{n=1}^{N}\tilde{\theta}_{n}^{t+1}.$
This aggregation ensures that the global model $\theta_{G}^{t+1}$ benefits from the knowledge of all clients while maintaining privacy through the DP mechanism.
5.3. Future Research Directions
Throughout this survey, we have extensively explored the applications of FL in SGs and identified vulnerabilities associated with FL-based SG systems. FL frameworks are commonly utilized across different stages of SGs, including generation, transmission/distribution, and consumption, with the primary goal of preserving privacy. However, there are significant gaps between SOTA FL methodologies and their practical implementation in SGs.
A major challenge is the issue of non-IID data among SG IoT devices. Since patterns vary significantly across devices, traditional FL approaches struggle to maintain model accuracy. Addressing non-IID data requires adaptive techniques that can handle diverse data distributions without compromising performance. Another area of focus is the deployment of large neural networks in FL, which can lead to computational inefficiencies and increased communication overhead. Optimizing the size and efficiency of models in FL contexts is crucial to scaling FL applications in SG environments. Additionally, improving aggregation strategies is essential to ensure effective collaboration between clients. Techniques such as weighted aggregation based on data quality and client reliability can reduce the noise and improve overall model accuracy. Client selection is also an important research direction—intelligent selection mechanisms that prioritize clients with high-quality data and sufficient computational resources can significantly enhance the efficiency and robustness of FL systems in SGs.
On the security front, addressing various types of attacks in FL-based SG systems is paramount. Byzantine attacks, for instance, pose a serious threat as malicious clients can send incorrect model updates, corrupting the global model. These attacks are particularly challenging because they blend in with honest client contributions, making them difficult to detect. Techniques such as robust aggregation methods (e.g., Krum and Bulyan) have been proposed, but further refinement is needed to protect FL systems from Byzantine failures. Backdoor attacks are another prominent threat in which adversaries implant hidden malicious behaviors into the global model while still achieving high performance on regular tasks. This is especially concerning in SG applications, where unnoticed tampering could lead to severe disruptions. Enhancing defense mechanisms such as differential privacy and anomaly detection algorithms can mitigate the risk of backdoor attacks. Inference attacks, where adversaries attempt to extract sensitive information from the shared model, are also a critical area for research. While FL inherently provides some protection by not sharing raw data, adversaries can still exploit the model’s updates to infer private details. Improving privacy-preserving techniques, such as homomorphic encryption and secure multi-party computation, can help protect against these inference attacks. To defend FL-based SG systems from these various threats, several advanced defense mechanisms are being developed. Byzantine-resilient aggregation methods like Trimmed-Mean (trmean_median), Krum (krum), and Bulyan (bulyan) offer robust solutions by filtering out malicious model updates. These approaches, however, need further exploration in the context of SGs, where the scale and complexity of the system introduce unique challenges. FedVRDP (zhang2023byzantine) is another promising approach that combines differential privacy with Byzantine resilience, offering a dual-layered defense against both privacy breaches and malicious attacks. Research into these and other defense strategies will be essential to making FL a viable and secure solution for real-world SG applications.
In conclusion, addressing these challenges and exploring new defense strategies will be critical in advancing FL-based SG systems. The integration of advanced aggregation strategies, efficient model optimization, and robust defense mechanisms will help close the gap between current research and practical applications, ensuring that FL can be securely and efficiently deployed in SG environments.
@@ 6. Conclusions
In this paper, we present a comprehensive review of studies focusing on FL-based applications in SGs and their potential vulnerabilities. We begin by comparing our work with existing surveys and outlining our unique contributions. Following this, we define key terminologies related to SGs and FL, providing a brief overview of commonly discussed attack methods within FL frameworks. Next, we examine the deployment of FL in three key stages of SG systems: generation, transmission and distribution, and consumption. The integration of FL at these stages has significantly enhanced data privacy and distribution efficiency in SGs. Furthermore, we analyze and compare the potential vulnerabilities inherent in FL-based SG systems. Finally, we summarize existing research on FL vulnerabilities and propose directions for future research.

### v3

@@ 4.2.2. SOTA Defenses
Detection-Filtering Methods. Detection-filtering methods aim to maximize robustness against attacks by identifying and removing malicious local model updates from the aggregation process. These methods typically rely on one or more statistical metrics (e.g., Cosine similarity and Euclidean distance) to assess local model updates. By comparing the extracted metrics with each update, they filter out updates that exhibit abnormal metric values, preventing them from influencing the global model. For example, Krum (Blanchard et al., 2017; Bhagoji et al., 2019) selects the most reliable local model that has the smallest sum of squared Euclidean distances to all other models as the output. Multi-Krum (Blanchard et al., 2017) extends this by selecting multiple reliable models for averaging instead of just one. Xu and Huang (2022) applied the KL-divergence to measure the distance between the empirical sign distributions of the local models in two consecutive rounds and used it to employ a weighted average of models for each client. The DnC proposed in (Shejwalkar and Houmansadr, 2021) calculates the product of a model with its principal component and removes the models with large projection values. FLTrust (Cao et al., 2020) leverages a root dataset and trains a root global model based on it. After collecting local updates, FLTrust aggregates those updates with high Cosine similarity to root model updates. SignGuard (Xu et al., 2022) combines direction-based clustering and magnitude-based filtering to identify malicious model updates, where the direction and magnitude are obtained by the sign statistics and $L_{2}$-norm of the model update, respectively. Recently, a defense method called LASA that achieves SOTA results has been proposed in (Xu et al., 2025b). It employs a layer-wise adaptive filter to adaptively select benign sparsified layers using both magnitude and direction metrics across all clients for aggregation.
@@
Impact-Reduction Methods. Impact-reduction methods try to integrate all model updates but employ strategies to reduce the impact of malicious updates. For instance, Trimmed mean (Trmean) proposed in (Yin et al., 2018a) discards a certain percentage of the highest and lowest values among the received models from clients for each dimension. After this trimming, the mean of the remaining values is computed by the server, which mitigates the impact of extreme values on the aggregated model. Guerraoui et al. (2018) proposed Bulyan that combines Trmean and Multi-Krum to extract benign models. Pillutla et al. (2022) proposed to use the geometric median (GeoMed) of all local models as the aggregated model, which is effective in filtering out the outlier values. RLR (Ozdayi et al., 2021) uses the sign information of model updates to adjust the global learning rate for each model coordinate without distinguishing between benign and malicious updates, resulting in poor accuracy. Foolsgold (Fung et al., 2020) assumes that the malicious updates are consistent with each other. It assigns aggregation weights to model updates based on the maximum Cosine similarity between the last layers of pairwise model updates. A higher Cosine similarity value indicates a higher probability that the updates are malicious, leading to smaller aggregation weights being assigned. SparseFed proposed in (Panda et al., 2022) sparsifies the aggregated model update at the server side, integrating with model clipping and error feedback, to mitigate the impact of malicious local model updates. In addition, FedVRDP proposed in (Zhang and Hu, 2023) applies sparsification on local model updates using a mask obtained from the global model in the previous round and combines it with random Gaussian noise perturbation and variance reduction techniques, to achieve robustness and differential privacy protection simultaneously. A recent approach called Lockdown (Huang et al., 2024) assumes that those malicious parameters served to recognize backdoor triggers will be deemed unimportant for benign clients and applies sparsification to prune unimportant parameters on the local model.
DP-based Methods. The goal of DP is to protect the participation of clients during the training of FL. Specifically, to achieve client-level DP, at training round $t$, each client $i$ clips their local model update $\Delta_{i}^{t}$ with a clipping threshold $C$. Subsequently, a small amount of DP noise, drawn from a Gaussian distribution $N(0,C^{2}\sigma^{2}/r\cdot\mathbf{I}_{d})$, is applied to the clipped model updates, where $\sigma^{2}$ is the noise multiplier and $r$ is the number of clients selected to participate in local training per round. When the server receives all perturbed model updates from clients, the accumulative noise becomes $N(0,C^{2}\sigma^{2}\cdot\mathbf{I}_{d})$, yielding a rigorous privacy guarantee.
However, adding noise to the model updates inherently degrades the model performance of an FL system. Therefore, many methods have been proposed to achieve better performance while maintaining privacy guarantees. For example, the first work to propose client-level DPFL is (McMahan et al., 2017b), which achieves client-level DP guarantees by employing the Gaussian mechanism and composing privacy guarantees. Cheng et al. (2022b) explicitly regularize the $L_{2}$-norm of local updates to ensure they are bounded and further mitigate the negative effects of clipping by applying post-training $\text{Top}_{k}$ local update sparsification. Fed-SMP (Hu et al., 2023a) assumes the server possesses a dataset similar to the overall distribution of client datasets. Using this server dataset, the server selects a set of $\text{Top}_{k}$ coordinates for all clients to prevent privacy leakage through these selected coordinates. AE-DPFL (Zhu et al., 2020) ensures user-level DP by employing a voting-based mechanism combined with secure aggregation. Unlike other approaches, AE-DPFL does not require clipping of the model or data, thereby mitigating accuracy degradation. However, AE-DPFL assumes the availability of unlabeled data from the global distribution on the server, which is often impractical in real-world applications. Shi et al. (2023) observe that DPFL methods tend to produce sharper loss landscapes, which can make models more sensitive to noise perturbations. To address this issue, they propose DP-FedSAM, which leverages the Sharpness-Aware Minimization (SAM) optimizer. This approach enhances parameter robustness against noise and aims to achieve more stable convergence points. Despite their promise, these methods often assume a uniform privacy requirement across all clients, which is impractical in real-world scenarios.
4.3. Discussions and Future Research Directions
Unique Challenges of FL-based Attack and Defense Methods in SGs. Deploying FL in SGs presents unique and substantial challenges that go beyond typical FL implementations, primarily due to the critical nature, scale, and inherent characteristics of these energy infrastructures. Successfully bridging the gap between SOTA FL methodologies and their practical, secure application in SGs requires addressing these multifaceted issues.
One of the foremost challenges stems from the resource-constrained nature of many SG devices. Components like smart meters, sensors, and actuators often possess limited computational power, memory, and energy budgets. This directly impacts the complexity of local model training permissible within an FL framework and constrains the sophistication of on-device operations. From an attack perspective, while simple disruptive actions like naive Byzantine attacks (e.g., sign-flips or random noise injection) might be feasible from such constrained devices, more advanced, computationally intensive attacks (e.g., optimization-based Byzantine attacks or complex local data poisoning for backdoors) might be limited to more powerful edge devices or orchestrated by a well-resourced adversary controlling multiple simpler nodes. Similarly, the implementation of robust on-device defenses or complex local differential privacy mechanisms can be severely hampered, forcing reliance on server-side or hierarchical defense strategies.
The integration of FL with existing SG infrastructure and its real-time operational demands poses another significant hurdle. SGs are often a hybrid of legacy systems and modern digital components, leading to interoperability issues, diverse communication protocols (each with potential vulnerabilities), and difficulties in retrofitting new technologies like FL securely. The real-time requirements for many SG applications, such as state estimation, fault detection, or dynamic load balancing, mean that FL models must be highly accurate and responsive. This critical dependency makes the SG system particularly vulnerable to attacks that degrade model performance or introduce latency. For example, Byzantine attacks that corrupt model updates or backdoor attacks that cause misbehavior on specific trigger conditions can have severe physical consequences, from economic losses to power outages. Consequently, defense mechanisms must not only be effective but also computationally efficient to operate within these low-latency constraints, and their false positive rates must be exceptionally low to prevent disruption of essential grid services.
Furthermore, SG data characteristics and governance complexities uniquely challenge FL deployment. SG data is inherently heterogeneous and non-IID due to varying consumer behaviors, geographical factors, and diverse energy resources. This non-IID nature can impede the convergence of global FL models, introduce biases, and significantly complicate the task of server-side defenses. Detection-filtering methods (e.g., Krum, statistical outlier removal) may struggle to distinguish between genuinely malicious updates and those that are merely statistically different due to legitimate local data variations. The sensitive nature of energy consumption data also makes SGs a prime target for inference attacks, such as Membership Inference Attacks, where adversaries attempt to glean private information from model updates or outputs. While FL aims to protect raw data, the model itself becomes a conduit for potential leakage. Implementing defenses like DP is crucial, but the inherent trade-off between privacy and model utility is particularly acute in SGs; the noise introduced to ensure privacy can degrade the accuracy of critical control and forecasting functions, demanding meticulous calibration.
Finally, the overall attack surface and the practical feasibility of robust defenses in a distributed SG environment present an overarching challenge. While FL decentralizes data, it introduces numerous potential points of failure or compromise through its participating clients (SG devices). Implementing sophisticated, coordinated attacks (e.g., ByzMean, distributed backdoors) becomes a question of an attacker’s ability to compromise a sufficient number of these distributed entities. On the defense side, there’s a constant tension between the need for robust, computationally intensive aggregation-stage defenses at the server (or edge aggregators) and the communication overhead and trust assumptions associated with collecting necessary information (like update similarities or auxiliary data) from potentially untrustworthy participants. The selection and tuning of defenses—whether detection-based, impact-reduction, or privacy-enhancing—must be carefully tailored to the specific SG application, the capabilities of participating devices, and the acceptable risk tolerance for critical infrastructure.
Future Directions for FL in Smart Grids. To advance the practical and secure deployment of FL in SG systems, future research must strategically address fundamental security, privacy, and operational challenges. A primary focus should be the development of highly tailored security and privacy frameworks robust enough for critical SG infrastructure. This involves creating adaptive, resource-aware defense mechanisms capable of mitigating sophisticated attacks (e.g., Byzantine, backdoor, inference) within heterogeneous SG environments characterized by resource-constrained devices and non-IID data. Concurrently, significant advancements are required in utility-preserving privacy-enhancing technologies, such as refined DP schemes or practical hybrid cryptographic solutions, to protect sensitive SG data during critical applications like state estimation or load forecasting, without unduly compromising model accuracy or operational efficiency, while also ensuring verifiable compliance with regulatory standards.
Furthermore, enhancing the operational viability and real-world applicability of FL in dynamic SG settings is paramount. A key direction is advancing asynchronous FL (AFL) protocols, which are inherently better suited to the SG’s typical device heterogeneity, variable network conditions, and potential for intermittent client connectivity than traditional synchronous approaches. Building upon existing research into Byzantine-resilient asynchronous methods (Damaskinos et al., 2018; Yang and Li, 2021; Fang et al., 2022), future work should develop lightweight yet robust AFL aggregation techniques specifically for SG data characteristics. This includes analyzing the distinct impact of asynchronicity on various attack surfaces and defense dynamics, ensuring model convergence and fairness under realistic SG conditions, and designing effective incentive mechanisms. Successfully tailoring and securing AFL for SGs will significantly improve system efficiency, scalability, and overall resilience in complex, large-scale deployments.
5. FedGridShield: An Open-Source Framework for FL-based Attack and Defense Methods in Smart Grids
We introduce FedGridShield [^1] https://github.com/FedGridShield/FedGridShield, a lightweight and scalable framework designed to implement FL-based attack and defense methods in SGs. FedGridShield supports ongoing cybersecurity research in the interdisciplinary field of FL and SGs by providing a user-friendly platform capable of handling various attack and defense methods and SG datasets.
The current framework includes attack methods such as AGR-tailored Trimmed-mean Attack (Shejwalkar and Houmansadr, 2021) and Fang Attack (Fang et al., 2020); defense methods including Trimmed Mean (Yin et al., 2018a), Median (Yin et al., 2018b), Krum (Blanchard et al., 2017), Bulyan (Guerraoui et al., 2018), and FLAME (Nguyen et al., 2022); and privacy-preserving methods DPFed (Zhang and Hu, 2023) and SparseFed (Panda et al., 2022).
We conduct experiments on two datasets: Electricity Theft Detection: We use data introduced by Zhao et al. (2023), which contains the power consumption data of 200 consumers, resulting in a total of 107,200 records. This is a tabular dataset with 48 features, and the binary label defines whether electricity theft is detected. Generator Defect Classification: We also collect a dataset of generator images in perfect and defective conditions. This dataset is designed to train an image classifier to indicate the generator’s condition, which can help with system maintenance. The training data contains 766 images of generators in perfect condition and 744 images of generators in defective condition. The test data contains 275 and 300 images for these conditions, respectively. The results are reported on the project page due to page restrictions.
@@ 6. Conclusions
In this paper, we present a comprehensive review of studies focusing on FL-based applications in SGs and their potential vulnerabilities. We begin by comparing our work with existing surveys and outlining our unique contributions. Following this, we define key terminologies related to SGs and FL, outline common FL threat models, and provide an overview of frequently discussed attack methods within FL frameworks. Next, we examine the deployment of FL across the three key stages of SG systems: generation, transmission and distribution, and consumption. The integration of FL at these stages has been shown to significantly enhance data privacy and distribution efficiency in SGs. Furthermore, we analyze the potential vulnerabilities inherent in these FL-based SG systems. We also introduce the FedGridShield framework, designed to facilitate further research and help address identified gaps. Finally, this paper summarizes existing research on FL vulnerabilities and proposes key directions for future investigation.
Acknowledgements.
The work of S. Rath was supported by the Faculty Startup Grant at the University of Tulsa.